OpenStudy (gudetamas):

A ladder sits 6 m away from the side of a house. The top of the latter rests at 14 m above the ground. The angle of elevation is 66.8^\circ . What is the length of the ladder?

OpenStudy (mathmale):

Please clarify: "The angle of elevation is 66.8^\circ"

OpenStudy (gudetamas):

ok

OpenStudy (gudetamas):

66.8 degrees i meant

OpenStudy (agent0smith):



OpenStudy (welshfella):

sin 66.8 = 14 / L solve for L

OpenStudy (gudetamas):

so i would do (L/1)66.8= 14/L (L/1) next to then get 14/sin(66.8)?

OpenStudy (welshfella):

yes